OpenStudy (anonymous):

A salesperson earns a salary of $10,000 per year. she also makes a commision of 26% on her sales. If her sales for the year total $21,000, how much money was she paid for the year?

OpenStudy (anonymous):

26% 0f 21000+10000

OpenStudy (anonymous):

\[10,000+.26\times 21,000=10,000+5460=15460\] check my arithmetic.

OpenStudy (anonymous):

you r right :)

OpenStudy (anonymous):

:)